If the magnitude of the vectors A,B and C are 6,8 and 10 units respectively and if A +B=C, then the angle between A and C is ?
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
B = C – A
 
on taking dot product to self
 
B^2 = C^2 + A^2 – 2 A.C 
 
64 = 100 + 36 – 2 * 6 * 10 * cos@
 
72 / 120 = cos@
 
cos @ = 6/10
 
cos@ = 3/5
 
@ = cos^-1 (3/5)